OpenStudy (anonymous):

The kinetic energy of an object moving at a speed of 16.8 m/s was determined to be 870 J. What is the mass of the object?

OpenStudy (anonymous):

Ek=1/2mv^2 870J=1/2m(16.8)^2 divide both sides by (16.8)^2 3.08=1/2m divide both sides by 1/2 6.16=m Mass= 6.16kg